A survey of 500 randomly chosen adults found that
20% of the 300 men and 15% of the 200 women had purchased books online.
You are to test the hypothesis that online rates are the same for men and women.
What is the p-value of your test of significance?
Answer #1

Answer

let p1 and p2 be the proportion of male and female

it is given that p1 = 0.20 with n1 = 300 and p2 = 0.15 with n2 = 200

p = (p1*n1 + p2*n2)/(n1+n2) = (0.2*300 + 0.15*200)/(500) = 0.18

z test statistic =

using z table, check 1.4 in the left most column and 0.03 in the first row, then select the intersecting cell for one tailed hypothesis, we get

p value = 0.0764